OpenStudy (anonymous):

What is the domain of f(x) = log(x + 3) + 1?

OpenStudy (lgbasallote):

will you draw how this looks like?

OpenStudy (anonymous):

I dont have anything that tells me wha it looks like? The only thing i have is this question and the answer options

OpenStudy (lgbasallote):

i meant what does log(x+3) + 1 look like...i need to see where the base and power are

OpenStudy (anonymous):

thats all i have for that problem nothing else...

OpenStudy (lgbasallote):

no...

OpenStudy (lgbasallote):

i just need to know how log(x+3) + 1 looks like...

OpenStudy (lgbasallote):

im not asking for exra info

OpenStudy (lgbasallote):

does it look like \[\huge \log (x+3) + 1\] or something else?

OpenStudy (anonymous):

OH! Im sooo sorry yes it sure does

OpenStudy (lgbasallote):

hmm here's a hint... if you have something like \[\huge \log a\] the domain is a>0 if you have something like \[\huge \log (x+a)\] the domain is x+ a> 0 then solve for x does that help? or do you need more help?

OpenStudy (anonymous):

more please...

OpenStudy (lgbasallote):

here's an example the domain of \[\huge \log x\] is x > 0 the domain of \[\huge \log (x-5)\] is x -5 > 0 or x > 5 the domain of \[\huge \log (x+2)\] is x + 2 > 0 or x > -2 does that help now?

OpenStudy (anonymous):

Yes! Very much thank you!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

What is the domain of f(x) = log(x + 4) - 2? ok well what about this one?

OpenStudy (lgbasallote):

it just follows the same thing

OpenStudy (anonymous):

ok....-4?

OpenStudy (anonymous):

im not the best at math as you can obviously see...sorry

OpenStudy (lgbasallote):

x > -4 dont forget x > it's very important

OpenStudy (anonymous):

oh yeah imost deffinately cant well can you stick around and help a bit more?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

What is the domain of f(x) = log(x + 2)?

OpenStudy (lgbasallote):

what do you think?

OpenStudy (anonymous):

x>-2?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Which of the following is the equation of the graph?

OpenStudy (anonymous):

hold on a sec... i need to attatch something

OpenStudy (anonymous):

what would i do for this?

OpenStudy (lgbasallote):

...what do you mean?

OpenStudy (anonymous):

sorry ive done that question already dont worry about it but i do need help with a few more...